genes A and B are 20 map units apart. If you cross Ab/aB xab/ab, how many each progeny type would you expect if total progeny was 1000.

a. AaBb(AB/ab)

b. Aabb ( Ab/ab)

c. aaBb(aB/ab)

d. aabb(ab/ab)

is b and c the parental ? would that mean that b and c are 20 ? Thanks
Answer #1

Answer:

Distance between genes (m.u.) = Recombination frequency (%)

Ab/aB genotype produces 4 types of gametes as below:

Recombinant gametes= 20%

AB = 10%

ab = 10%

Non-recombinant gametes = 80%

Ab = 40%

aB = 40%

ab/ab genotype produces only one type of gametes, ab.

Ab/aB xab/ab –Parents

Ab

Ab(40%)

Ab/ab(40%)

aB(40%)

aB/ab (40%)

AB (10%)

AB/ab (10%)

ab(10%)

ab/ab (10%)

a. AaBb(AB/ab) = 10/100 * 1000 = 100

b. Aabb ( Ab/ab) = 40/100 * 1000 = 400

c. aaBb(aB/ab) = 40/100 * 1000 = 400

d. aabb(ab/ab) = 10/100 * 1000 = 100

The b & c are parental.